### v1

# Fluctuation-dissipation and virtual processes in interacting phonon systems

Abstract
Phonon-phonon interactions are fundamental to understanding a wide range of material properties, including thermal transport, expansion and vibrational spectra. In all of the conventional perturbative approaches, energy conservation is enforced using Dirac delta functions at each microscopic phonon interaction. We demonstrate below that these delta functions stem from an incomplete treatment of the interacting problem, that violates the fluctuation-dissipation theorem governing systems at equilibrium. By replacing the deltas with convolutions of the phonon spectral functions, and introducing a self-consistency condition, we provide a more accurate and physically consistent framework. For systems where phonon dynamics can be approximated as Markovian, we simplify this approach, reducing the dissipative component to a single parameter tied to phonon lifetimes. Applying this method to boron arsenide, we find that self-consistent linewidths capture the phonon scattering processes better, significantly improving agreement with experimental thermal conductivity values. These results also challenge the conventional view of four-phonon processes as dominant in BAs, demonstrating a three-phonon description can explain most of the experimental value, provided it is self-consistent. With this method we address critical limitations of perturbative approaches, offering new insights into dissipation and phonon-mediated processes, and enabling more accurate modeling of anharmonic materials.
preprint: APS/123-QED
Phonons are quantized vibrational modes of a crystal lattice, and central to our understanding of a wide range of material properties. From mechanical stability and sound propagation to optical responses and thermal transport, phonons play a key role in determining the behavior of solids, and enable many functionalities and applications. As such, an accurate description of phonon dynamics is essential to advance our understanding of both fundamental physics and material functionality. In the harmonic approximation, phonons are treated as non-interacting quasiparticles with infinite lifetimes. This can naturally not be a complete description: in real systems, anharmonicity introduces phonon-phonon interactions, which lead to a broadening of phonon spectral lineshapes. These interactions are crucial to describe key material properties quantitatively and sometimes even qualitatively. For example, without interactions the phonon thermal conductivity is infinite, an inconsistency which is resolved with the introduction of anharmonicity.
In conventional perturbative approaches, phonon-phonon interactions are treated with strict energy conservation, enforced by Dirac delta functions of the energy difference before and after each scattering event [1, 2, 3, 4]. In this letter, we show that while this formalism captures the idealized dynamics of non-interacting phonons, it employs a strong approximation that neglects the dissipative component of real phonon dynamics, and breaks the fluctuation-dissipation theorem. We show that a more exact formulation replaces delta functions with spectral function convolutions and a self-consistency condition, which can be interpreted as high-order effective interactions mediated by virtual phonons. Importantly, these are not subject to energy conservation. A practical implementation is introduced for cases where phonon dynamics are approximated as Markovian, reducing the dissipative component to a single parameter associated with phonon lifetimes.
We are interested in phonon correlation functions at finite temperature, which are linked to observables such as X-ray or neutron scattering. For the derivation, we place ourselves within our recently introduced mode-coupling theory of anharmonic lattice dynamics [5, 6, 7], but similar arguments will hold for all of the commonly used (perturbative) theories. Within this formalism, the phonon Kubo correlation functions (KCF) [8] obey the generalized Langevin equation of motion
[…]
where we introduce the super-index $\lambda=(\mathbf{q},s)$, with $\mathbf{q}$ the wave vector in reciprocal space and $s$ the index of the phonon branch. This equation describes the dynamics as driven by a harmonic force with the spring magnitude given by the phonon frequency $\Omega_{\lambda}$, and a time-dependent friction term. The latter is contained in a memory kernel $K_{\lambda}(t)$ which describes the dissipation of the phonon energy through a bath made up of all other phonons. Importantly, $K_{\lambda}(t)$ must respect the fluctuation-dissipation theorem [9, 10, 8], which states that $K_{\lambda}(t)=\big{(}\delta f_{\lambda},\delta f_{\lambda}(t)\big{)}$, where $\delta f_{\lambda}(t)$ is the “random” force which describes the effects of neglected degrees of freedom on the dynamics. The dynamical susceptibility (or spectral function) is obtained through a Laplace transform of eq.(1), giving
$\chi_{\lambda}^{\prime\prime}(\omega)=\frac{1}{\hbar\pi}\frac{4\omega\Omega_{\lambda}\Gamma_{\lambda}^{2}(\omega)}{(\omega^{2}-\Omega_{\lambda}^{2}-2\omega\Delta_{\lambda}(\omega))^{2}+4\omega^{2}\Gamma_{\lambda}^{2}(\omega)},$ (2)
where $\Gamma_{\lambda}(\omega)$ and $\Delta_{\lambda}(\omega)$ are respectively the real and imaginary parts of its memory kernel (which are related by a Kramers-Kronig transform). The main complication of eq.(2) lies in the expression of the memory kernel. Within the mode-coupling approximation, $K$ is obtained by projecting the component of the “random” force onto multiple displacements, introducing interactions between phonons. To simplify our presentation, we will stay at the lowest order of phonon-phonon interaction, but it is important to note that our reasoning holds identically with higher orders. With only three-phonon interactions and after a decoupling of four-point correlations into products of two-point correlations [7], the real part of the memory kernel has the form
$\Gamma_{\lambda}(\omega)=\frac{1}{16}\sum_{\lambda^{\prime}\lambda^{\prime\prime}}|\Psi_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}|^{2}S_{\lambda^{\prime}\lambda^{\prime\prime}}(\omega).$ (3)
In this equation, $\Psi_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}$ are third-order scattering matrix elements, and $S$ is a function describing the effect of phonons $\lambda^{\prime}$ and $\lambda^{\prime\prime}$ on the memory kernel. $S$ is expressed as
[…]
where the lesser phonon correlation function $G_{\lambda}^{<}(\omega)$ is proportional to the KCF, and related to the dynamical susceptibility by the fluctuation-dissipation theorem, through $G_{\lambda}^{<}(\omega)=n(\omega)\chi_{\lambda}^{\prime\prime}(\omega)$. $n=(e^{\beta\hbar\omega}-1)^{-1}$ is the Bose-Einstein distribution.
A common simplification involves approximating $S_{\lambda^{\prime},\lambda^{\prime\prime}}(\omega)$ using the memory-less dynamical susceptibility of the phonons $\lambda^{\prime}$ and $\lambda^{\prime\prime}$. This results in $S_{\lambda^{\prime},\lambda^{\prime\prime}}(\omega)$ being expressed as a sum of delta functions centered at sums or differences of phonon frequencies, weighted by phonon occupation numbers [7]
$\begin{split}S_{\lambda^{\prime}\lambda^{\prime\prime}}(\omega)&=(n_{\lambda^{\prime}}+n_{\lambda^{\prime\prime}}+1)\delta(\omega-\Omega_{\lambda^{\prime}}-\Omega_{\lambda^{\prime\prime}})\\ &-(n_{\lambda^{\prime}}+n_{\lambda^{\prime\prime}}+1)\delta(\omega+\Omega_{\lambda^{\prime}}+\Omega_{\lambda^{\prime\prime}})\\ &+(n_{\lambda^{\prime}}-n_{\lambda^{\prime\prime}})\delta(\omega+\Omega_{\lambda^{\prime}}-\Omega_{\lambda^{\prime\prime}})\\ &-(n_{\lambda^{\prime}}-n_{\lambda^{\prime\prime}})\delta(\omega-\Omega_{\lambda^{\prime}}+\Omega_{\lambda^{\prime\prime}}),\end{split}$ (5)
where $n_{\lambda}=n(\Omega_{\lambda})$. The delta functions in this equation have a clear physical interpretation: they represent energy conservation during each phonon interaction. However, this approach neglects a critical element: the memory kernel. In a memory-less framework, phonons are idealized as infinitely long-lived quasiparticles, with their dynamical susceptibility given by $\chi_{\lambda}^{\prime\prime}(\omega)\propto\delta(\omega-\Omega_{\lambda})$. This idealization inherently assumes non-dissipative dynamics, so that energy conservation arises automatically from the system’s behavior. Introducing the memory kernel fundamentally changes this picture. Phonons are no longer viewed as perfectly conserving energy through their dynamics; instead, the memory kernel accounts for the exchange of energy between interacting phonons. In this framework, the energy conservation for individual events is generalized to the detailed balance condition, which gives rise to the fluctuation-dissipation theorem. Energy conservation should only be found between the initial and final states. Consequently, the delta functions, which cannot capture dissipation, are replaced by a convolution term in eq.(3), which is self-consistently linked to eq.(2). This convolution reflects the interplay between phonons and their self-created environment, where the memory kernel describes the “phonon bath” formed by the collective dynamics of all phonons. Modifying the memory kernel alters the phonon properties, which in turn reshape the bath—a feedback mechanism captured by the self-consistent framework.
Figure 1: Examples of second-order three-phonon Feynman diagrams for anharmonic systems. The black wiggly lines represent phonon propagators, the red ones representing intermediate virtual phonons and the blue dots represent 3-phonon vertices. a. Second order decay process. b. Second order redistribution process.
In a diagrammatic framework, the convolution can be interpreted as introducing higher-order interactions mediated by virtual phonons [11, 12]. For instance, the process illustrated in FIG.1a. represents an effective four-phonon interaction, where one phonon decays into three phonons through an intermediate process involving a virtual phonon. Similarly, FIG.1b. depicts a second-order redistribution process in which two phonons merge to form a virtual phonon, which subsequently decays into two new phonons. In such higher-order interactions, energy and momentum conservation are strictly enforced only for the initial and final phonons. The virtual phonons, however, are not subject to these constraints, enabling phonon interactions that would otherwise be forbidden by strict energy conservation. When self-consistency is applied, even more complex higher-order interactions emerge, and at full convergence, an infinite series of such interactions contributes to the phonon linewidths. Consequently, these previously forbidden interactions, which are simply a manifestation of the fluctuation-dissipation condition, can have drastic consequences on the properties of systems where selection rules could restrict anharmonicity [13, 14, 15, 16, 17]. For instance, in systems with large mass differences - and thus a large acoustic-optical band gap - interactions between acoustic and optical modes are no longer forbidden within a three-phonon truncation. This should modify substantially the scattering phase space and reduce, for example, the expected large conductivity in these systems.
The self consistency of the full frequency dependent dynamical susceptibility represents an important computational challenge due to the convolution appearing in $S_{\lambda^{\prime},\lambda^{\prime\prime}}(\omega)$. However, phonons are often accurately described using a Markovian approximation, where the memory kernel reduces to a single number $\Gamma_{\lambda}=\Gamma_{\lambda}(\Omega_{\lambda})$ and $\chi_{\lambda}^{\prime\prime}(\omega)$ to a sum of Lorentzian centered on ±$\pm\Omega_{\lambda}$ with widths $\Gamma_{\lambda}$. In this case, the $S_{\lambda^{\prime},\lambda^{\prime\prime}}(\omega)$ function has the simple form [7]
$\begin{split}S_{\lambda^{\prime}\lambda^{\prime\prime}}(\omega)&=(n_{\lambda^{\prime}}+n_{\lambda^{\prime\prime}}+1)L(\omega,\Omega_{\lambda^{\prime}}+\Omega_{\lambda^{\prime\prime}},\Gamma_{\lambda^{\prime}}+\Gamma_{\lambda^{\prime\prime}})\\ &-(n_{\lambda^{\prime}}+n_{\lambda^{\prime\prime}}+1)L(\omega,-\Omega_{\lambda^{\prime}}-\Omega_{\lambda^{\prime\prime}},\Gamma_{\lambda^{\prime}}+\Gamma_{\lambda^{\prime\prime}})\\ &+(n_{\lambda^{\prime}}-n_{\lambda^{\prime\prime}})L(\omega,\Omega_{\lambda^{\prime}}-\Omega_{\lambda^{\prime\prime}},\Gamma_{\lambda^{\prime}}+\Gamma_{\lambda^{\prime\prime}})\\ &-(n_{\lambda^{\prime}}-n_{\lambda^{\prime\prime}})L(\omega,-\Omega_{\lambda^{\prime}}+\Omega_{\lambda^{\prime\prime}},\Gamma_{\lambda^{\prime}}+\Gamma_{\lambda^{\prime\prime}}),\end{split}$ (6)
where $L$ are Lorentzian functions defined as $L(\omega,\Omega,\Gamma)=\Gamma/((\omega-\Omega)^{2}+\Gamma^{2})/\pi$. Thus, within the Markovian approximation, the fluctuation-dissipation theorem can be simply enforced by using this equation in a self-consistent cycle involving equations (2), (3) and (6). This represents a simplification compared to the numerical implementation of delta functions [18, 2, 3, 4] as it avoids the need of introducing numerical convergence parameters and usually creates a broader smearing than would be used numerically, easing convergence. It does however bring an extra-cost related to the need to compute eq.(3) several times. It should be noted that the need for linewidth self-consistency has already been proposed in previous work [19, 20]. However, these works do not propose a formal justification for this self-consistency. Moreover, the smearing used in these works does not correspond to the convolution of the Lorentzian dynamical susceptibility appearing in eq.(6), breaking the relation between the dissipative component of the phonon dynamics and the memory kernel, and with it, the fluctuation-dissipation theorem.
To demonstrate the importance of the fluctuation-dissipation condition, we applied our approach to boron arsenide, a semiconductor renowned for its exceptionally high thermal conductivity [21, 22, 23, 24]. A key feature of this material is its large band gap separating the acoustic and optical phonon branches, as shown in FIG.2a. This gap is sufficiently wide to forbid three-phonon coupling between acoustic and optical phonons if strict energy conservation is applied [15], making four-phonon interactions a requirement to accurately reproduce the thermal conductivity in this system [25]. However, when the energy conservation principle is replaced by the fluctuation-dissipation theorem, third-order acoustic-optical interactions are no longer strictly forbidden: through the Markovian approximation physically natural Lorentzian functions [7] appear. Their long tails extend the scattering phase space, significantly altering the dynamics of both acoustic and optical phonons. This highlights BAs as an archetypal system where the fluctuation-dissipation theorem should play a critical role in capturing the true phonon scattering processes, and their impact on thermal transport.
Our results in FIG.2 were computed using machine-learning interatomic potentials constructed on ab initio data, whose details are provided in the SM [7].
Figure 2: a. T-dependent phonons in BAs from 100 to 1100 K. The large acoustic to optical band-gap is highlighted in grey. b. Thermal conductivity in BAs including 3 or 3+4 phonon interactions, with energy conservation enforced or using the fluctuation-dissipation condition. Experimental results from Kang et al [21], Li et al [22] and Tian et al [23] are shown for comparison. c. Comparison of the phonon linewidths at $300$ K with 3-phonon interactions with energy conservation vs fluctuation-dissipation condition. d. Comparison of the spectral thermal conductivity at $300$ K computed with energy conservation (ec) or after self-consistency to enforce the fluctuation-dissipation condition (fd). The phonon density of states is represented in grey.
At the three-phonon level, the impact of the fluctuation-dissipation condition is particularly striking. While enforcing strict energy conservation allows us to reproduce theoretical results from the literature [24, 25, 15], incorporating self-consistent linewidths consistently leads to lower thermal conductivity values, which are in closer agreement with experimental measurements [23, 21, 22]. This reduction can be attributed to a significant decrease in the thermal conductivity contribution from phonons with energies around $20$ meV. Under strict energy conservation, these phonons are forbidden from interacting with optical modes, but the convolution in the self-consistent approach relaxes this restriction, thereby expanding the scattering phase space. This behavior is analogous to what occurs when four-phonon processes are included: the increased scattering phase space results in higher linewidths for these phonons and a corresponding reduction in their thermal conductivity contribution. Our findings challenge the prevailing view that four-phonon interactions are essential to explain the thermal conductivity in large phonon band gap materials like boron arsenide. While these interactions are not negligible, their importance appears to be overestimated. Instead, remaining at the three-phonon level while enforcing the fluctuation-dissipation condition provides a reasonably accurate prediction of thermal transport in this material, consistent with its low anharmonicity.
Although the fluctuation-dissipation condition is universally applicable, its impact on different materials varies, and requires quantitative evaluation. In the SM [7], we demonstrate that the impact of self-consistency is negligible for silicon — a highly harmonic system with low linewidths and no phonon band gap — whereas it is significant for silver iodide, which exhibits strong anharmonicity. Ultimately, we anticipate that self-consistency will substantially alter the phonon dynamics in systems with large mass differences or pronounced anharmonicity. For example, hydrogen-rich materials, which feature both characteristics, are promising candidates for further exploration of the fluctuation-dissipation condition’s role in phonon interactions.
While our study focuses on phonons, similar self-consistent approaches should be relevant for other finite-temperature quasiparticles. Systems involving interactions between quasiparticles with vastly different energy scales, such as magnon-phonon or electron-phonon interactions, will be particularly affected by the relaxation of strict energy conservation. For instance, recent work [26, 27] demonstrates that incorporating self-consistent electronic linewidths resolves divergences previously expected in piezoelectric materials. This highlights the broader importance of capturing dissipative dynamics in a rigorous and consistent framework, extending the relevance of our approach beyond phonons to other complex quasiparticle systems.
Acknowledgements.
We thank S. Poncé and J-M. Lihm for discussions about self-consistent electron phonon calculations. The authors acknowledge the Fonds de la Recherche Scientifique (FRS-FNRS Belgium) and Fonds Wetenschappelijk Onderzoek (FWO Belgium) for EOS project CONNECT (G.A. 40007563), and Fédération Wallonie Bruxelles and ULiege for funding ARC project DREAMS (G.A. 21/25-11). MJV acknowledges funding by the Dutch Gravitation program “Materials for the Quantum Age” (QuMat, reg number 024.005.006), financed by the Dutch Ministry of Education, Culture and Science (OCW). Simulation time was awarded by by PRACE on Discoverer at SofiaTech in Bulgaria (optospin project id. 2020225411); EuroHPC-JU award EHPC-EXT-2023E02-050 on MareNostrum 5 at Barcelona Supercomputing Center (BSC), Spain; by the CECI (FRS-FNRS Belgium Grant No. 2.5020.11); and by the Lucia Tier-1 of the Fédération Wallonie-Bruxelles (Walloon Region grant agreement No. 1117545).

### v2

@@
 Our findings challenge the prevailing view that four-phonon interactions are essential to explain the thermal conductivity in large phonon band gap materials like boron arsenide.
-While these interactions are not negligible, their importance appears to be overestimated.
-Instead, remaining at the three-phonon level while enforcing the fluctuation-dissipation condition provides a reasonably accurate prediction of thermal transport in this material, consistent with its low anharmonicity.
-Although the fluctuation-dissipation condition is universally applicable, its impact on different materials varies, and requires quantitative evaluation.
-In the SM [7], we demonstrate that the impact of self-consistency is negligible for silicon — a highly harmonic system with low linewidths and no phonon band gap — whereas it is significant for silver iodide, which exhibits strong anharmonicity.
-Ultimately, we anticipate that self-consistency will substantially alter the phonon dynamics in systems with large mass differences or pronounced anharmonicity.
-For example, hydrogen-rich materials, which feature both characteristics, are promising candidates for further exploration of the fluctuation-dissipation condition’s role in phonon interactions.
+While these interactions are not negligible, their importance appears to have been overestimated in previous literature.
+Remaining at the three-phonon level while enforcing the fluctuation-dissipation condition provides a reasonably accurate prediction of thermal transport in this material, consistent with its low anharmonicity.
+However, this accuracy should not be seen as an indication that the self-consistent determination of linewidth could avoid the need for four-phonon interactions.
+Instead, the fluctuation-dissipation condition should be seen as a new, superior framework to evaluate all scattering mechanisms.
+Its importance in the context of BAs stems from the constrained 3-phonon scattering phase-space produced by strict energy conservation.
+The FD approach removes this constraint, and shows the reproduction of the experimental thermal conductivity of BAs using energy conservation and the 4-phonon channel is actually a coincidence: the EC 3-phonon $\kappa$ is too high, and the EC 4-phonon correction is too large, but happens to land near the experimental values.
+The fluctuation-dissipation condition is universally applicable, but its impact on different materials varies, and requires quantitative evaluation.
+In the SM [7], we show the impact of the self-consistency for silicon, a highly harmonic system with low linewidths and no phonon band gap, and for silver iodide, which exhibits strong anharmonicity.
+In the former case, the difference in using the fluctuation-dissipation condition is negligible, both with three and four-phonon interactions.
+In AgI however, its impact is more pronounced.
+Including only three-phonon scattering, the self-consistency reduces the thermal conductivity by 15% to 25% between 100 to 300 K.
+When four phonon scattering is included, the impact of the fluctuation-dissipation condition is smaller, between 5% and 20 % in the same range of temperature, which is still significant for low thermal conductivity materials.
+Ultimately, the fluctuation-dissipation condition is a universal principle that applies to all materials, yet its impact on phonon dynamics appears to be highly system-dependent.
+Our findings highlight the need for broader investigations to determine how this condition interacts with anharmonicity and the available scattering phase space.
+In more complex or strongly interacting systems, additional factors, such as possible deviations from Markovian behavior, may also influence the outcome.
+Such studies will be essential for establishing when and how fluctuation-dissipation-consistent treatments are necessary for accurate modeling of lattice dynamics.
 While our study focuses on phonons, similar self-consistent approaches should be relevant for other finite-temperature quasiparticles.